LaToya throws a ball from the top of a bridge. Her throw is modeled by the equation y=-.5x^2+3x+12 and the bridge is modeled by the equation y=-.2x+9. About how far does the ball travel horizontally before its first​ bounce?
1
Expert's answer
2020-11-15T18:15:44-0500

"y_1=-5x^2+3x+12",

"y_2=-2x+9",

"y_1=y_2",

"5x^2-5x-3=0",

"x_1=\frac{1-\sqrt{\frac{17}{5}}}{2}" ,

"x_2=\frac{1+\sqrt{\frac{17}{5}}}{2}" ,

"x_2-x_1=\sqrt{\frac{17}{5}}\approx1.84."